A Spherical water tank is to be constructed out of plastic materials. If the radius of the tank is 2 meters , and thickness of the water tank is to be 0.5 dm. What is the volume of the materials needed to construct the tank?


1
Expert's answer
2021-06-21T13:45:37-0400

"V=\frac{4}{3}\pi*1^3-\frac{4}{3}\pi*(1-0.05)^3=0.5974 m^3."